# Learning Actionable World Models for Industrial Process Control

### v2

@@ I Introduction
Automatic process monitoring plays a central role in maintaining repeatable and stable processes with high-quality products in industrial processes [1]. Commonly, the monitoring task is formulated as an anomaly detection task, as it is a priori unknown how exactly a process may deviate from normal operations [2]. To then bring it back to its intended operational step once an anomaly is detected, it is necessary to identify and remove the cause of any detected anomaly. Existing approaches focus on anomaly detection [3] and lack interpretability, making it difficult to identify the underlying root causes. One reason is that real-world monitored processes—such as plastic injection molding or air condition flow—often lack an analytical or simulation model of the relationship between measured process signals and control parameters (actions: physical signals, such as mold temperature adjustments, directly applied to the system, e.g., an injection molding machine, to regulate its behavior). This complicates root cause analysis for control despite reliable algorithms for anomaly detection. Thus, understanding relations between process signals and actions is crucial to transition from passive monitoring to active control.
@@
In this paper, we introduce a novel world model for process control, inspired by joint embedding predictive architectures (JEPAs) and contrastive learning [6, 9]. Our world model makes use of recorded process signals and corresponding actions to learn an action-aware latent representation, facilitating disentangled action prediction in the presence of many possible control parameter adjustments. Experimental results on a real-world use case from plastic injection molding control demonstrate that our world model achieves its goals with high sample efficiency (limited industrial data of only 80 samples). Our main contributions are: (a) a world model for process control, based on a novel architecture, inspired by the JEPA framework and contrastive learning, conditions the learned latent representation of the world’s state on the difference in parameters evoking this state as compared to a reference state; (b) experimental validation on a real-world injection molding case, where the contrastive modeling of pressure curve pairs and corresponding machine parameters allows the model to better distinguish between subtle variations in the data, showing disentanglement for different actions in the embedding space for the first time as compared to the state of the art; (c) a novel evaluation scheme for the quality of disentanglement, based on angle and distance deviations between ground truth and predicted actions and showing the validity of our approach for real-world process control.
@@ II Background and related work
Joint embedding representation learning. The quality of input reconstruction from a generative model does not necessarily correspond to the quality of the learned representations [24]. JEPAs have emerged as a promising SSL approach for representation learning, offering an alternative to generative models: Instead of taking a loss in the reconstructed space, they predict and compute the loss in latent space, conditioned on actions. Recent self-supervised variants (I-JEPA [25] and V-JEPA [26]) predict masked parts of the input, learning representations for computer vision tasks. We extend the JEPA framework by moving away from an implicit masking action ($\boldsymbol{x}\xrightarrow{\boldsymbol{masking}}\boldsymbol{y}$) to explicitly providing arbitrary actions in a factored form ($\boldsymbol{x}\xrightarrow{\boldsymbol{a}}\boldsymbol{y}$), facilitating learning a rich latent representation of cause and effect economically. Thus, we extend the typical SSL framework by not being agnostic to external factors (like machine parameters).
@@
Disentangled representation learning for time-series. Deep learning has been proposed to learn effective representations of the time-series data for downstream applications [27, 28, 12, 29]. To be useful for fine-grained process control, the learned representations must be disentangled: A low-dimensional latent representation $\boldsymbol{z}=\{z_{1},z_{2},\ldots,z_{n}\}$ is sought such that all pairs of latent variables in $Z$ are independent, i.e., the change of the sequential patterns in the input corresponding to variable $z_{i}$ is invariant to variable $z_{j}$, denoted as $z_{i}\perp\!\!\!\perp z_{j}$, $\forall i,j$, where $i,j\in\{1,2,\ldots,n\}$ and $i\neq j$. The disentanglement can be extended to group disentanglement so that a group of latent variables is independent of other different groups. Most representation learning relies on unsupervised learning due to the label scarcity. However, unsupervised learning of disentangled representations is fundamentally impossible without inductive biases on both the models and the data [30]. While interpretable semantic concepts often depend on the interaction of multiple factors rather than individual components, Li et al. argue that it is essential to interpret time-series data through a single latent representation in industrial applications [31]. A disentangled representation can be decomposed into separate dimensions, where the change in each dimension reflects a distinctive change in the real world. Li et al. propose to learn interpretable time-series representations through encouraging high mutual information between the latent codes and the original time-series to achieve disentangled latent representations [31]. Hamaguchi et al. [14] use contrastive learning to learn common and specific features during representation learning. Then, only the common features are used for downstream tasks. In contrast, we address disentanglement through action-conditioned world modeling.
@@ III Methodology
Figure 2: Proposed actionable world model architecture (general case; inputs for injection molding example shown, see Sec. IV). There are $3$ main components in the model, encoder $Enc$, latent predictor $P_{z}$, and action predictor $P_{a}$. Losses are taken in representation space and for the action prediction.
@@ IV-B Datasets
Figure 3: Illustration of the three-dimensional machine parameter space of D1/D2 with a normalized scale. The gray points represent $27$ different machine parameters settings. The green and red points represent the most extreme settings in the corners of the cube (outline indicated by blue lines for clarity), in which individual parameters take on either minimal or maximal values. The pressure curves shown in the dashed rectangles show three examples of a reference signal $\boldsymbol{x}$ (green) and observed signal $\boldsymbol{y}$ (red) associated machine parameter setting; the arrows represent an action to transition from $\boldsymbol{x}\xrightarrow{\boldsymbol{a}}\boldsymbol{y}$.
@@
Data preprocessing. The pressure time-series represent the dynamic changes in pressure over time, which are critical for understanding the process’s behavior. As the original time-series data is sampled with irregular time intervals, we preprocess the data by resampling the data to a standardized format using $500$ time points with regular intervals between 00 and $10$ seconds. This interpolation ensures uniformity across all pressure curve samples. Additionally, we normalize pressure data by dividing by $1000$, roughly rescaling the observed data to a range between 00 and $1$.

### v3

@@ I Introduction
Automatic process monitoring plays a central role in maintaining repeatable and stable processes with high-quality products in industrial processes [1]. Commonly, the monitoring task is formulated as an anomaly detection task, as it is a priori unknown how exactly a process may deviate from normal operations [2]. To restore the system to its normal operational status once an anomaly is detected, it is necessary to identify and remove the cause of any detected anomaly. Existing approaches focus on anomaly detection [3] and lack interpretability, making it difficult to identify the underlying root causes. One reason is that real-world monitored processes—such as plastic injection molding or air condition flow—often lack an analytical or simulation model of the relationship between measured process signals and control parameters (actions: physical signals, such as mold temperature adjustments, directly applied to the system, e.g., an injection molding machine, to regulate its behavior). This complicates root cause analysis for control despite reliable algorithms for anomaly detection. Thus, understanding relations between process signals and actions is crucial to transition from passive monitoring to active control.
@@
In this paper, we introduce a novel world model for process control, inspired by joint embedding predictive architectures (JEPAs) and contrastive learning [6, 9]. Our world model makes use of recorded process signals and corresponding actions to learn an action-aware latent representation, facilitating disentangled action prediction in the presence of many possible control parameter adjustments. Experimental results on a real-world use case from plastic injection molding control demonstrate that our world model achieves its goals with high sample efficiency (limited industrial data of only 80 samples). Our main contributions are: (a) a world model for process control, based on a novel architecture, inspired by the JEPA framework and contrastive learning; it conditions the learned latent representation of the world’s state on the difference in parameters evoking this state as compared to a reference state; (b) experimental validation on a real-world injection molding case, where the contrastive modeling of pressure curve pairs and corresponding machine parameters allows the model to better distinguish between subtle variations in the data, showing disentanglement for different actions in the embedding space for the first time as compared to the state of the art; (c) a novel evaluation scheme for the quality of disentanglement, based on angle and distance deviations between ground truth and predicted actions and showing the validity of our approach for real-world process control.
@@ II Background and related work
Joint embedding representation learning. The quality of input reconstruction from a generative model does not necessarily correspond to the quality of the learned representations [24]. JEPAs have emerged as a promising SSL approach for representation learning, offering an alternative to generative models: Instead of taking a loss in the reconstructed space, they predict and compute the loss in latent space, conditioned on a latent variable (e.g., actions) [6]. Recent JEPA variants (I-JEPA [25] and V-JEPA [26]) predict masked parts of the input, learning representations for computer vision tasks. We extend the JEPA framework by moving away from an implicit masking action ($\boldsymbol{x}\xrightarrow{\boldsymbol{masking}}\boldsymbol{y}$) to explicitly providing arbitrary actions in a factored form ($\boldsymbol{x}\xrightarrow{\boldsymbol{a}}\boldsymbol{y}$), facilitating learning a rich latent representation of cause and effect economically. Thus, we extend the typical SSL framework by not being agnostic to external factors (like machine parameters).
@@
Disentangled representation learning for time-series. Deep learning has been proposed to learn effective representations of the time-series data for downstream applications [27, 28, 12, 29]. To be useful for fine-grained process control, the learned representations must be disentangled: A low-dimensional latent representation $\boldsymbol{z}=\{z_{1},z_{2},\ldots,z_{n}\}$ is sought such that all pairs of latent variables in $Z$ are independent, i.e., the change of the sequential patterns in the input corresponding to variable $z_{i}$ is invariant to variable $z_{j}$, denoted as $z_{i}\perp\!\!\!\perp z_{j}$, $\forall i,j$, where $i,j\in\{1,2,\ldots,n\}$ and $i\neq j$. The disentanglement can be extended to group disentanglement so that a group of latent variables is independent of other different groups. Most representation learning relies on unsupervised learning due to label scarcity. However, unsupervised learning of disentangled representations is fundamentally impossible without inductive biases on both the models and the data [30]. While interpretable semantic concepts often depend on the interaction of multiple factors rather than individual components, Li et al. argue that it is essential to interpret time-series data through a single latent representation in industrial applications [31]. A disentangled representation can be decomposed into separate dimensions, where the change in each dimension reflects a distinctive change in the real world. Li et al. propose to learn interpretable time-series representations through encouraging high mutual information between the latent codes and the original time-series to achieve disentangled latent representations [31]. Hamaguchi et al. [14] use contrastive learning to learn common and specific features during representation learning. Then, only the common features are used for downstream tasks. In contrast, we address disentanglement through action-conditioned world modeling.
@@ III Methodology
Figure 2: Proposed actionable world model architecture (general case; inputs for injection molding example shown, see Sec. IV). There are $3$ main components in the model, encoder $Enc$, latent predictor $P_{z}$, and action predictor $P_{a}$. Losses are computed in representation space and for the action prediction.
@@ IV-B Datasets
Figure 3: Illustration of the three-dimensional machine parameter space of D1/D2 with a normalized scale. The gray points represent $27$ different machine parameters settings. The green and red points represent the most extreme settings in the corners of the cube (outline indicated by blue lines for clarity), in which individual parameters take on either minimal or maximal values. The pressure curves shown in the dashed rectangles show one example of a reference signal $\boldsymbol{x}$ (green) and two observed signal $\boldsymbol{y}$ (red), each corresponding to different machine parameter settings; the arrows represent an action to transition from $\boldsymbol{x}\xrightarrow{\boldsymbol{a}}\boldsymbol{y}$.
@@
Data preprocessing. The pressure time-series represent the dynamic changes in pressure over time, which are critical for understanding the process’s behavior. As the original time-series data is sampled with irregular time intervals, we preprocess the data by resampling the data uniformly to $500$ time points between 00 and $10$ seconds. This interpolation ensures uniformity across all pressure curve samples. Additionally, we normalize pressure data by dividing by $1000$, roughly rescaling the observed data to a range between 00 and $1$.